# A Corpus for Named Entity Recognition in Chinese Novels with Multi-genres

### v1

@@ Abstract
Entities like person, location, organization are important for literary text analysis. The lack of annotated data hinders the progress of named entity recognition (NER) in literary domain. To promote the research of literary NER, we build the largest multi-genre literary NER corpus containing 263,135 entities in 105,851 sentences from 260 online Chinese novels spanning 13 different genres. Based on the corpus, we investigate characteristics of entities from different genres. We propose several baseline NER models and conduct cross-genre and cross-domain experiments. Experimental results show that genre difference significantly impact NER performance though not as much as domain difference like literary domain and news domain. Compared with NER in news domain, literary NER still needs much improvement and the Out-of-Vocabulary (OOV) problem is more challenging due to the high variety of entities in literary works.
@@ 2 Related Work
Currently, there is relatively little research on NER in the literary field due to the diverse types of entities and significant differences in naming styles and background knowledge(Labatut and Bost, 2019). Establishing a general NER model for the literary field is challenging, and the lack of large-scale NER datasets limits the development of NER research in this domain(Augenstein et al., 2017). Several previous studies have proposed different approaches and built different corpora for named entity recognition in literary works. Vala et al. (2015) introduce a graph-based pipeline model specifically for character recognition. Brooke et al. (2016) propose the LitNER model, which utilizes the bootstrap method. Bamman et al. (2019) build LitBank corpus, annotating named entities in 100 English novels and trained an NER model tailored to the literary field. Dekker et al. (2019) conduct an evaluation of natural language processing tools to extract characters and build social networks from novels. For Chinese literary NER, Xu et al. (2017) construct a dataset for NER and relationship extraction from essays. In addition, we Jia et al. (2021) create a named entity dataset using Jin Yong’s novels and develop a named entity recognition model that incorporates document-level information. The overview statistics of above datasets are shown in Table 1.
@@ 3.2 Inter-annotator Agreement
To ensure consistent and high-quality annotation, we adopt a multi-round iterative approach. Two annotators simultaneously annotate each novel, cross-check and review each other’s work, guaranteeing reliable results. The annotation process consisted of two stages: experimental and formal annotation. In the experimental stage, we use the LTP Che et al. (2020) named entity recognition tool to pre-annotate the novels’ text, gain familiarity with the corpus and improve the annotation guidelines. In the formal annotation stage, one annotator initially annotates the text, which is then verified by a second annotator to resolve any inconsistencies. The final results are confirmed by the first annotator. This process involves seven annotators and is completed in 70 days. We assess annotation consistency using the F1 score as the evaluation metric. Results show a micro-averaged F1 score (MicroF1) of 92.15% and a Micro-averaged F1 score (MacroF1) of 88.62%, indicating high reliability of the dataset Artstein and Poesio (2008). The consistency varies across entity types, with person entities demonstrating higher consistency compared to organization and location entities. The complex structures of organization and location entities pose challenges in identifying their boundaries. Detailed values are given in Table 2.
@@ 3.3 Corpus Analysis
Genre Type
High-frequency entities
PER
球迷(Fan),教练 (Coach),于指导(Guidance Yu ),职业球员(Professional player)
Sport LOC
中国(China),美国(USA),英格兰(England),西雅图(Seattle)
ORG
NBA,青年队(youth teams),森林队(forest teams)
PER
太宰(Taizai),崇祯(Chongzhen),大魏天子(Emperor of Wei),刘总管(General Manager Liu)
History LOC
秦国(Qin State),京城(Capital),汴梁(Bianliang),宜城(Yicheng)
ORG
锦衣卫(Jinyiwei),中书省(Zhongshu Province),东宫卫队(the Eastern Palace Guard),豫山书院(Yushan Academy)
PER
老法师(Old Mage),女巫(Witch),黑衣武士(Black-clad Warrior),魔法师(Magician)
Fantasy LOC
城堡(Castle),鲜花镇(Flower Town),乌山镇(Wushan Town),荆棘岭(Thornridge),圣域(Sanctum)
ORG
神盾局(S.H.I.E.L.D.),魔法学院(Academy of Magic),死局帮(Deadlock Gang)
PER
玄幽道人(Xuanyou Taoist),独孤败天(Dugu Baitian),刘三刀(Liu Sandao),司徒傲月(Situ Aoyue)
Wuxia LOC
蜀山(Shu Mountain),华山(Hua Mountain),通州(Tongzhou),中原(Central Plains)
ORG
飞鹰帮(Feiying Gang),画剑派(Huajian Sect),李家(Li Family)
PER
青年导游(Youth Tour Guide),赢胖子(Fatty Ying),王秘书(Secretary Wang),副经理(Deputy Manager)
Urban LOC
商场(Shopping malls),网吧(internet cafes),中云市(Zhongyun City),办公室(offices),江南(Jiangnan)
ORG
医院(hospitals),学府(academic institutions),大学(universities),战争学府(war academies)
@@ 4.5 Case Study
Ground Truth(from Game)
这位女伯爵${}_{PER}$$\cdots$都被视为匈牙利${}_{LOC}$最美的女人之一。
BiLSTM-CRF
这位女伯爵${}_{PER}$$\cdots$都被视为匈牙利${}_{PER}$最美的女人${}_{PER}$之一。
BERT-CRF
这位女伯爵${}_{PER}$$\cdots$都被视为匈牙利${}_{PER}$最美的女人${}_{PER}$之一。
BERT-BiLSTM-CRF
这位女伯爵${}_{PER}$$\cdots$都被视为匈牙利${}_{PER}$最美的女人${}_{PER}$之一。
Ground Truth(from Xuanhuan)
古河${}_{PER}$名列加玛帝国${}_{ORG}$十大强者之一。
BiLSTM-CRF
古河${}_{PER}$名列加玛帝国${}_{ORG}$十大强者之一。
BERT-CRF
古河${}_{LOC}$名列加玛帝国${}_{ORG}$十大强者之一。
BERT-BiLSTM-CRF
古河${}_{LOC}$名列加玛帝国${}_{ORG}$十大强者之一。

### v2

@@ Abstract
Entities like person, location, organization are important for literary text analysis. The lack of annotated data hinders the progress of named entity recognition (NER) in literary domain. To promote the research of literary NER, we build the largest multi-genre literary NER corpus containing 263,135 entities in 105,851 sentences from 260 online Chinese novels spanning 13 different genres. Based on the corpus, we investigate characteristics of entities from different genres. We propose several baseline NER models and conduct cross-genre and cross-domain experiments. Experimental results show that genre difference significantly impact NER performance though not as much as domain difference like literary domain and news domain. Compared with NER in news domain, literary NER still needs much improvement and the Out-of-Vocabulary (OOV) problem is more challenging due to the high variety of entities in literary works. Our data and models are open-sourced at https://github.com/hjzhao73/MultiGenre-ChineseNovel.
@@ 2 Related Work
Currently, there is relatively little research on NER in the literary field due to the diverse types of entities and significant differences in naming styles and background knowledge(Labatut and Bost, 2019). Establishing a general NER model for the literary field is challenging, and the lack of large-scale NER datasets limits the development of NER research in this domain(Augenstein et al., 2017).
Several previous studies have proposed different approaches and built different corpora for named entity recognition in literary works. Vala et al. (2015) introduce a graph-based pipeline model specifically for character recognition. Brooke et al. (2016) propose the LitNER model, which utilizes the bootstrap method. Bamman et al. (2019) build LitBank corpus, annotating named entities in 100 English novels and trained an NER model tailored to the literary field. Dekker et al. (2019) conduct an evaluation of natural language processing tools to extract characters and build social networks from novels. For Chinese literary NER, Xu et al. (2017) construct a dataset for NER and relationship extraction from essays. In addition, we Jia et al. (2021) create a named entity dataset using Jin Yong’s novels and develop a named entity recognition model that incorporates document-level information. The overview statistics of above datasets are shown in Table 1.
@@ 3.2 Inter-annotator Agreement
To ensure consistent and high-quality annotation, we adopt a multi-round iterative approach. Two annotators simultaneously annotate each novel, cross-check and review each other’s work, guaranteeing reliable results. The annotation process consisted of two stages: experimental and formal annotation. In the experimental stage, we use the LTP Che et al. (2020) named entity recognition tool to pre-annotate the novels’ text, gain familiarity with the corpus and improve the annotation guidelines. In the formal annotation stage, one annotator initially annotates the text, which is then verified by a second annotator to resolve any inconsistencies. The final results are confirmed by the first annotator. This process involves seven annotators and is completed in 70 days.
We assess annotation consistency using the F1 score as the evaluation metric. Results show a micro-averaged F1 score (MicroF1) of 92.15% and a Micro-averaged F1 score (MacroF1) of 88.62%, indicating high reliability of the dataset Artstein and Poesio (2008). The consistency varies across entity types, with person entities demonstrating higher consistency compared to organization and location entities. The complex structures of organization and location entities pose challenges in identifying their boundaries. Detailed values are given in Table 2.
@@ 3.3 Corpus Analysis
Genre Type High-frequency entities
PER 球迷(Fan),教练 (Coach),于指导(Guidance Yu ),职业球员(Professional player)
Sport LOC 中国(China),美国(USA),英格兰(England),西雅图(Seattle)
ORG NBA,青年队(youth teams),森林队(forest teams)
PER 太宰(Taizai),崇祯(Chongzhen),大魏天子(Emperor of Wei),刘总管(General Manager Liu)
History LOC 秦国(Qin State),京城(Capital),汴梁(Bianliang),宜城(Yicheng)
ORG 锦衣卫(Jinyiwei),中书省(Zhongshu Province),东宫卫队(the Eastern Palace Guard),豫山书院(Yushan Academy)
PER 老法师(Old Mage),女巫(Witch),黑衣武士(Black-clad Warrior),魔法师(Magician)
Fantasy LOC 城堡(Castle),鲜花镇(Flower Town),乌山镇(Wushan Town),荆棘岭(Thornridge),圣域(Sanctum)
ORG 神盾局(S.H.I.E.L.D.),魔法学院(Academy of Magic),死局帮(Deadlock Gang)
PER 玄幽道人(Xuanyou Taoist),独孤败天(Dugu Baitian),刘三刀(Liu Sandao),司徒傲月(Situ Aoyue)
Wuxia LOC 蜀山(Shu Mountain),华山(Hua Mountain),通州(Tongzhou),中原(Central Plains)
ORG 飞鹰帮(Feiying Gang),画剑派(Huajian Sect),李家(Li Family)
PER 青年导游(Youth Tour Guide),赢胖子(Fatty Ying),王秘书(Secretary Wang),副经理(Deputy Manager)
Urban LOC 商场(Shopping malls),网吧(internet cafes),中云市(Zhongyun City),办公室(offices),江南(Jiangnan)
ORG 医院(hospitals),学府(academic institutions),大学(universities),战争学府(war academies)
@@ 4.5 Case Study
[…]
BiLSTM-CRF $\cdots她都被视为匈牙利PER$最美的女人PER之一。
BERT-CRF $\cdots她都被视为匈牙利PER$最美的女人PER之一。
BERT-BiLSTM-CRF $\cdots她都被视为匈牙利PER$最美的女人PER之一。
Ground Truth(from Xuanhuan) 古河PER名列加玛帝国ORG十大强者之一。
BiLSTM-CRF 古河PER名列加玛帝国ORG十大强者之一。
BERT-CRF 古河LOC名列加玛帝国ORG十大强者之一。
BERT-BiLSTM-CRF 古河LOC名列加玛帝国ORG十大强者之一。